# Jailbreak Attacks and Defenses against Multimodal Generative Models: A Survey

### v1

@@ A.1 Input-level Defense
Input analysis typically involves examining statistical features of the input, including detecting sensitive keywords, and assessing perplexity levels.
Statistics-based Defense. Statistical-based methods typically rely on established statistical metrics and rule-based criteria rather than complex models, making them easier to implement and interpret in various practical scenarios.
@@
$\bullet$ Text → Image (T2I) Models. In commercial T2I products such as Midjourney [31], Leonardo.ai [96] and DALL-E [32], blacklist methods are a commonly employed strategy for ensuring content safety and compliance with platform guidelines. These methods involve creating a predefined list of restricted or prohibited words and phrases that the T2I model must filter out or modify in its inputs.
@@ A.2 Encoder-level Defense
Encoder-level defense focuses on the examination of the latent space within the encoder of generative models to assess the semantic tendencies of latent representations.
Embedding-based Defense. Without relying solely on explicit keywords or rules, embedding-based defenses assess the semantics and similarities between embeddings.
@@
$\bullet$ Text → Image (T2I) Models. Latent Guard [98] detects unsafe T2I prompts by mapping the latent representations of blacklisted concepts and corresponding prompts within a shared latent space. The approach begins by utilizing a pre-trained text encoder to extract embeddings, followed by training an Embedding Mapping Layer that emphasizes relevant tokens through cross-attention mechanisms. Similarly, GuardT2I [99] optimizes a generative framework to convert latent embeddings into natural language, accurately capturing the user’s intent. By revealing the true intent of input prompts, GuardT2I then applies both blacklist and Sentence Similarity Checker to detect malicious prompts.
@@ A.3 Generator-level Defense
Generator-level defense involves distilling internal embeddings or analyzing the probability distributions to identify anomalous patterns indicative of harmful content.
Embedding-based Defense. These defenses leverage embedding values from the hidden states of the generative model to monitor suspicious activity.
$\bullet$ Image + Text → Text (IT2T) Models. NEARSIDE [100] introduces a method to detect adversarial inputs by identifying the attack direction within the hidden space of VLMs. This attack direction is derived by calculating the average embedding difference between benign and adversarial inputs. An input is then classified as adversarial if the projection of its embedding onto the obtained attack direction exceeds a defined threshold.
@@ A.4 Output-level Defense
Output-level Defenses typically employ iterative querying and specialized toxic detectors to assess the final outputs of the model.
Query-based Defense. Query-based methods involve repeatedly querying the model with perturbed inputs and analyzing the variance in the resulting outputs.
@@
Detector-based Defense. To enhance detection accuracy, an alternative approach is to train a dedicated toxicity classifier.
@@ IV-B Transformative Defense
Despite the improved accuracy in discriminative detection methods, these approaches often incorrectly flag benign inputs as harmful. Consequently, there is a growing research focus on transformative defenses that can operate at four levels to influence the model’s generation process, ensuring benign responses even in the presence of adversarial or malicious prompts in Fig. 8.
@@ B.1 Input-level Defense
Defenders can enhance input prompts by embedding defense cues at the input level, which act as safeguards to steer the model towards generating benign outputs when confronted with malicious prompts.
Prefix-based Defense. Prefix-based Defense entails appending carefully crafted prefixes to input prompts as a preventative measure, guiding the model toward generating benign content.
$\bullet$ Image + Text → Text (IT2T) Models. AdaShield [103] optimizes a prompt generator to construct a diverse pool of defense prompts tailored to various scenarios. During inference, this approach selects the most suitable defense prompt from the pool based on semantic similarity, appending it as a prefix to model inputs to guard against jailbreak attacks. Moreover, BlueSuffix [103] comprises three key components: a suffix generator fine-tuned through reinforcement learning, a textual purifier, and a visual purifier. Specifically, the suffix generator produces a defensive suffix, which is appended to the purified text input generated by the textual purifier. These modified text inputs, combined with the purified images from the visual purifier, are then used as defense inputs for the target VLMs. Unlike approaches that use proxy models to generate prefixes, UNIGUARD [105] applies gradient-based optimization to construct specialized safety guardrails for each modality. These guardrails including visual noises and textual suffixes, are optimized to minimize the likelihood of generating harmful responses in a toxic corpus.
@@
$\bullet$ Text → Image (T2I) Models. POSI [106] uses Proximal Policy Optimization to fine-tune a language model as an optimizer, transforming toxic prompts into clean ones. It introduces a novel reward function that assesses both the toxicity level and the textual alignment of the generated images.
@@ B.2 Encoder-level Defense
For transformative defense, encoder-level methods focus on modifying the encoder’s behavior and can be categorized into two types: fine-tuning during the training phase and guidance implementation during the inference phase.
@@
$\bullet$ Image + Text → Text (IT2T) Models. Sim-CLIP series methods [107, 108] introduce an unsupervised adversarial fine-tuning approach for the CLIP vision encoder to enhance robustness against adversarial attacks. The approach utilizes a siamese architecture, optimizing adversarial perturbations by maximizing cosine similarity between clean and perturbed image representations.
$\bullet$ Text → Image (T2I) Models. AdvUnlearn [109] proposes a utility-retaining regularization to optimize the trade-off between concept erasure robustness and model utility. It identifies the text encoder as more effective than UNet for robust unlearning and demonstrates its effectiveness in erasing nudity, objects, and style concepts across various diffusion models. Safe-CLIP [110] fine-tunes the CLIP model on synthetic data obtained from a LLM. It enhances safety in cross-modal tasks, including image-to-text retrieval, text-to-image generation, and image generation.
Guidance-based Defense. Guidance-based methods at the encoder level aim to embed specific information into the latent embeddings associated with potentially harmful content, steering the model’s behavior during content generation.
$\bullet$ Text → Image (T2I) Models. Self-discovery[111] proposes to operate on prompts implicitly. This approach discovers vectors representing desired concepts without requiring labeled data or external models. These vectors can be added to the embeddings of the condition prompt to guide responsible generation, promoting fairness and safety.
@@ B.3 Generator-level Defense
Generative-level defense involves editing the model during training and incorporating guidance or pruning mechanisms during inference, thereby directly influencing the model’s internal generation process.
Tuning-based Defense. These approaches involve modifying the model’s behavior during the training time through adjustments to parameters or optimizing objectives, such as revising harmful prompts, erasing unsafe concepts, or strategically employing safety modules.
$\bullet$ Image + Text → Text (IT2T) Models. A group of meth- ods [112, 113, 114, 115] primarily fine-tune models with supervised instruction datasets to ensure the model’s output probability distribution aligns more closely with desired safe outputs. This is formally expressed as:
$\mathop{\mathrm{arg}\min}\limits_{\theta}\;-\sum_{i=1}^{m}{\log\left(p_{\theta}\left(\mathcal{Y}_{i}|\mathcal{X}_{harm}\right)\right)}.$ (9)
Specifically, VL-Guard [112] presents a vision-language safety dataset covering both harmful and safe images paired with relevant instructions. This dataset is used to evaluate and fine-tune models with two types of methods: post-hoc fine-tuning and mixed fine-tuning. Apart from solely training MLLMs, SafeVLM [113] introduces additional safety modules including a safety projector, safety tokens, and a safety head. Through a two-stage training process, this approach progressively trains safety modules and fine-tunes the language model. Similarly, BaThe [114] employs a series of trainable soft embeddings as a “wedge” within MLLMs. This approach maps harmful instructions to rejection responses during training, effectively transforming the model into a defense mechanism against jailbreak attacks.
@@
$\bullet$ Text → Image (T2I) Models. Fine-tuning pre-trained T2I models to erasure unsafe concepts has received wide attention. To erase the concept, the model is trained to predict the negatively guided noises that move the model’s prediction away from the erased concept.
@@
$\mathop{\mathrm{arg}\min}\limits_{\epsilon_{\theta^{\prime}}}\mathbb{E}_{x_{t},t}\left[\lVert\epsilon_{\theta^{\prime}}(x_{t},e_{c},t)-\epsilon_{E}\rVert^{2}\right]$ (10)
$x_{t}\in\mathbb{R}^{W\times H\times d}$ is the denoised image at timestep $t$. $e_{c}$ is the text embedding for concept $c$. $\epsilon_{E}$ is the negatively guided noise predicted by $\theta$, $\epsilon_{E}=\epsilon_{\theta}(x_{t},t)-\eta\left[\epsilon_{\theta}(x_{t},e_{c},t)-\epsilon_{\theta}(x_{t},t)\right]$. By minimizing the distance between $\epsilon_{\theta^{\prime}}(x_{t},e_{c},t)$ and $\epsilon_{E}$, the model learns to lower the probability of the generated image $x$ aligning with the target concept $c$, thereby removing the undesired concept.
Addressing the trade-off between erasing unsafe concepts and preserving locality is a key problem to be solved. Receler [116] trains a lightweight eraser module to erase concepts while ensuring robustness and locality. This is achieved through a concept-localized regularization technique combined with an adversarial prompt learning scheme. EraseDiff [117] frames the unlearning task as a constrained optimization problem, balancing the removal of information related to the data to be forgotten while preserving the model’s performance on the retained data. It accomplishes this by modifying the generative process to diverge from the original denoising procedure associated with the forgotten data. ShieldDiff [130] tackles unsafe content removal in T2I models by fine-tuning a pre-trained diffusion model through reinforcement learning. It utilizes a content-safe reward function specifically designed to produce high-quality, safe images. This function integrates the CLIP model with nudity-specific rewards, enabling effective filtering of inappropriate content while preserving the intended semantics.
Some methods [118, 119, 120, 121] focus on the parameters of the attention module. Forget-Me-Not [118] introduces a plug-and-play method for efficiently forgetting and correcting specific concepts in large-scale text-to-image models. Attention Re-steering is designed to fine-tune only the UNet component to minimize the attention maps corresponding to the target concepts. SalUn [119] introduces the concept of “weight saliency”. By focusing on specific model weights rather than the entire model, SalUn enhances the efficiency and effectiveness of the unlearning process. UCE [120] addresses bias, copyright, and offensive content in text-to-image models by applying edits without additional training. It uses a closed-form solution to modify the model’s attention weights, selectively adjusting the keys and values associated with specific text embeddings for the concepts to be edited while minimizing alterations to the preserved concepts. MACE [121] expands the erasure capacity to handle up to 100 concepts by combining closed-form cross-attention refinement with LoRA fine-tuning, enabling efficient and scalable concept removal.
Several methods [122, 123, 124, 125, 126, 127] use the safe concept and the unsafe concept to guide the model. ESD [122] leverages a brief text description of the unsafe concept to guide the model. It fine-tunes the model by using conditioned and unconditioned scores from the frozen model, steering the output away from the concept being erased. This allows the model to use its internal knowledge to guide the diffusion process, effectively avoiding the undesired concept. SDD [123] applies self-distillation to the diffusion model, guiding the noise estimate conditioned on the target concept to align with the unconditional estimate. It enables the simultaneous removal of multiple concepts in a single process. SafeGen [124] focuses on eliminating explicit unsafe visual representations within the model, regardless of the text input. This approach makes the text-to-image model resistant to adversarial prompts, as it blocks the generation of such unsafe visuals from the inside. SFD [125] encourages the forgetting of undesirable information in diffusion models by aligning the conditional scores of “unsafe” classes or concepts with those of “safe” ones. To avoid relying on real data, the SFD framework integrates a score-based loss into the score distillation objective of a pre-trained diffusion model. UC [126] proposes a domain correction framework using adversarial training to align sensitive and anchor concepts in diffusion models, along with gradient surgery to preserve model performance by mitigating conflicts between unlearning and relearning gradients. Direct Unlearning Optimization (DUO) [127] removes unsafe content from T2I models through preference optimization with paired image data, while output-preserving regularization ensures the model maintains its generative capabilities on safe content.
Different from the previous methods, RECE [128] employs closed-form parameter editing combined with adversarial learning schemes to achieve reliable and efficient concept erasing, ensuring the effective removal of targeted concepts from the model. Another work, Dark Miner [129], follows a recurring three-stage process of mining, verifying, and circumventing. It greedily mines embeddings with the highest generation probabilities for unsafe concepts and effectively reduces the generation of unsafe content by targeting and mitigating these high-risk embeddings.
@@
$\bullet$ Image + Text → Text (IT2T) Models. InferAligner[131] proposes to align poorly aligned models for harmlessness by extracting safety steering vectors from pre-aligned models. It targets harmful inputs through intent recognition via safety related vector and utilizes a guidance gate to control activation shifts during inference. By selectively shifting activations in specific transformer layers, InferAligner ensures that only harmful intents are addressed, preserving the model’s performance on other tasks.
$\bullet$ Text → Image (T2I) Models. EIUP[132] introduces an erasure prompt that fuses image features with text embeddings. Through attention mechanisms, it identifies and re-weights non-compliant features, suppressing the generation of unsafe images from implicit unsafe prompts. SLD [93] mitigates inappropriate content generation by combining text conditioning with classifier-free guidance, removing or suppressing unsafe concepts in image outputs. It edits images during inference without fine-tuning, using “unsafe” prompts to guide generation in the opposite direction. Geom-Erasing [133] extracts geometric information of implicit concepts from images, converts it into location tokens, and incorporates these into text prompts. During sampling, these tokens serve as negative prompts, steering the model to generate images free from implicit concepts.
$\bullet$ Text → Video (T2V) Models. SAFREE [134] is a training-free framework that ensures safe T2I and T2V generation by identifying and redirecting prompt token embeddings away from toxic subspaces without altering model weights. It employs a self-validating filtering mechanism to balance toxicity filtering and the preservation of safe concepts, dynamically adjusting denoising steps and using adaptive re-attention within the diffusion latent space. This integration across both textual and visual spaces enables SAFREE to maintain output fidelity, quality, and safety across diverse generative models.
Pruning-based Defense. Unsafe content often originates from concept neurons, which are key to generating specific concepts. Pruning model parameters to deactivate unwanted concept neurons offers an effective strategy for eliminating harmful content.
P-ESD [135] enhances concept-erasing techniques by identifying concept-correlated neurons that are sensitive to adversarial prompts. It selectively removes critical parameters linked to the target concepts to remove undesired concepts. ConceptPrune [136] identifies skilled neurons in feed-forward layers responsible for undesirable concepts in diffusion models.

### v2

@@ A.1 Input-level Defense
Input analysis typically relies on established statistical metrics and rule-based criteria, making them easy to implement and interpret in various practical scenarios.
Statistics-based Defense. Examining statistical features of the input, including assessing perplexity levels and detecting sensitive keywords, is a common and efficient strategy.
@@
$\bullet$ Text → Image (T2I) Models. In commercial T2I products [31, 103, 32], blacklist methods are a commonly employed strategy for ensuring content safety and compliance with platform guidelines. These methods involve creating a predefined list of restricted or prohibited words and phrases that the T2I models must filter out in their inputs.
@@ A.2 Encoder-level Defense
Without relying solely on explicit keywords or rules, Encoder-level defense focuses on the examination of the latent space within the pre-trained encoders.
Perturbation-based Defense. Perturbation-based Defense within the encoder level involves disturbing inputs to observe the resulting variations in the generated embeddings.
@@
Proxy-based Defense. Proxy-based defenses pass the generated embeddings to the external models for analyzing the potential semantic tendencies.
$\bullet$ Text → Image (T2I) Models. Latent Guard [105] detects unsafe prompts by mapping the latent representations of blacklisted concepts and corresponding prompts within a shared latent space. The approach begins by utilizing a pre-trained text encoder to extract embeddings, followed by training an Embedding Mapping Layer that emphasizes relevant tokens through cross-attention mechanisms. Similarly, GuardT2I [106] optimizes a generative framework to convert latent embeddings into natural language, accurately capturing the user’s intent. By revealing the true intent of input prompts, GuardT2I then applies both blacklist and Sentence Similarity Checker to detect malicious prompts.
@@ A.3 Generator-level Defense
Generator-level defense involves distilling activations from internal hidden states to identify anomalous patterns indicative of harmful content.
Perturbation-based Defense. Apart from embeddings from the latent space, embedding values from the hidden states of generators can also be utilized to monitor suspicious activity.
$\bullet$ Image + Text → Text (IT2T) Models. NEARSIDE [107] identifies the attack direction from the hidden states by calculating the average embedding difference between benign and adversarial inputs. Inputs are then classified as adversarial if the projection of their embedding onto the obtained attack direction exceeds a defined threshold.
@@ A.4 Output-level Defense
Output-level Defenses typically employ iterative querying with perturbed inputs or specialized toxic detectors to assess the model outputs.
Perturbation-based Defense. Without relying on intermediate model values, an alternative approach is to analyze the variance in the resulting outputs by iteratively querying models with perturbed inputs.
@@
Proxy-based Defense. Training a dedicated external classifier to detect the toxicity of generated responses is widely used as an automatic evaluation method [109, 148, 149, 150, 151].
@@ IV-B Transformative Defense
Despite the improved accuracy in discriminative defense, these methods often incorrectly flag benign inputs as harmful. Recent research increasingly focuses on transformative defenses that can operate at four levels to influence the model’s generation process, ensuring benign responses when confronted with adversarial or malicious prompts, as shown in Fig. 8.
@@ B.1 Input-level Defense
Defenders can enhance inputs by embedding defense cues or applying purification via refiners to steer the model toward generating benign outputs.
Perturbation-based Defense. Appending carefully crafted prefixes or noises to inputs is viewed as an effective preventative measure.
$\bullet$ Image + Text → Text (IT2T) Models. AdaShield [110] employs a prompt generator to construct a diverse pool of defense prompts tailored to various scenarios. During inference, this approach selects the most suitable defense prompt from the pool based on semantic similarity, appending it as a prefix to model inputs. Unlike using proxy models to generate prefixes, UNIGUARD [111] applies gradient-based optimization to construct specialized safety guardrails for each modality. These guardrails including visual noises and textual suffixes, are optimized to minimize the likelihood of generating harmful responses in a toxic corpus.
@@
$\bullet$ Image + Text → Text (IT2T) Models. BlueSuffix [112] comprises three primary components: a textual purifier, a visual purifier, and a suffix generator. Specifically, the textual purifier rewrites adversarial prompts and appends the purified text with defensive suffixes generated by the suffix generator. These modified textual inputs, combined with the purified images from the visual purifier, serve as defense inputs.
$\bullet$ Text → Image (T2I) Models. POSI [113] uses proximal policy optimization to fine-tune a language model that transforms toxic prompts into clean ones. It introduces a novel reward function that assesses both the toxicity level and the textual alignment of the generated images.
@@ B.2 Encoder-level Defense
For transformative defense, encoder-level methods focus on modifying the encoder’s behavior and can be categorized into three types: fine-tuning during the training phase, applying guidance and refiner during the inference phase.
@@
$\bullet$ Image + Text → Text (IT2T) Models. Sim-CLIP series methods [59, 60] introduce an unsupervised approach to fine-tune the built-in vision encoder, aiming to enhance robustness against adversarial attacks. Specifically, the approach introduces a Siamese architecture combined with adversarial training strategies that maximize cosine similarity between clean and perturbed image representations.
$\bullet$ Text → Image (T2I) Models. Both AdvUnlearn [114] and Safe-CLIP [61] fine-tune the text encoder in T2I models. AdvUnlearn [114] proposes a utility-retaining regularization to optimize the trade-off between concept erasure robustness and model utility. It effectively erases nudity, objects, and style concepts across various diffusion models. Safe-CLIP [61] fine-tunes on synthetic data obtained from a LLM. It enhances safety in cross-modal tasks, including image-to-text retrieval, text-to-image generation, and image generation.
Guidance-based Defense. Embedding specific information into the latent space aims to deviate from the generation directions of unsafe content.
$\bullet$ Text → Image (T2I) Models. Self-discovery [115] proposes to add a learned latent vector to the embeddings of the condition prompt to guide responsible generation. These latent vectors are discovered without requiring labeled data or external models. Specifically, it first generates images from text prompts containing the target concept. These images are then denoised using a frozen diffusion model, guided by a modified prompt excluding the concept and an introduced learnable latent vector. By minimizing the reconstruction loss, the latent vector learns to represent the concept.
Refiner-based Defense. Since prompt embeddings are likely the source of unsafe generation, refiners can purify inappropriate concepts from embeddings into benign representations.
$\bullet$ Text → Image (T2I) Models. ES [116] introduces a plug-and-play module that operates on prompt embeddings. By using an internal scoring network that assigns harmfulness scores to individual tokens, ES identifies inappropriate concepts within prompt embeddings, enabling adaptive sanitization that prioritizes high-risk tokens while preserving benign ones.
@@ B.3 Generator-level Defense
Generative-level transformation involves editing the model during training and incorporating guidance or pruning mechanisms during inference, thereby directly influencing the model’s internal generation process.
Tuning-based Defense. These defenses modify the model’s behavior by adjusting parameters based on optimized objectives, such as likelihood loss and noise prediction loss.
$\bullet$ Image + Text → Text (IT2T) Models. Some meth- ods [117, 118, 119, 120] primarily fine-tune models with supervised instruction datasets to ensure the model’s output probability distribution aligns more closely with desired safe outputs $\left\{y_{i}^{s}\right\}_{i=1}^{m}$. This is formally expressed as:
$\mathop{\mathrm{arg}\min}\limits_{\theta}\;-\sum_{i=1}^{m}{\log\left(p_{\theta}\left(y_{i}^{s}|x_{mal}\right)\right)}.$ (10)
where $m$ is the sequence length and $p_{\theta}\left(\cdot\right)$ represents the probability of next token prediction. Specifically, VL-Guard [117] presents a vision-language safety dataset covering both harmful and safe images paired with relevant instructions, which are used to evaluate and fine-tune models. In addition to direct tuning, other works [118, 119] introduce additional safety modules. SafeVLM [118] incorporates a safety projector, safety tokens, and a safety head. Through a two-stage training process, this approach progressively trains safety modules and fine-tunes the language model. Similarly, BaThe [119] employs a series of trainable soft embeddings, aiming to map harmful instructions to rejection responses during training.
@@
$\bullet$ Text → Image (T2I) Models. To erase the concept, a group of works [122, 125, 124, 121, 126, 127, 128, 123] learn to lower the probability of the generated image aligning with the target concept $c$, which can be formulated as:
@@
$\mathop{\mathrm{arg}\min}\limits_{\theta}\mathbb{E}_{z_{t},t}\left[\lVert\epsilon_{\theta}(z_{t},e_{c},t)-\epsilon_{E}\rVert^{2}\right],$ (11)
where $z_{t}$ is the denoised image at timestep $t$. $\epsilon_{E}$ is the negatively guided noise. Specifically, ESD [121] and Receler [122] fine-tune the model by using conditioned and unconditioned scores from the frozen model, steering the output away from the concept being erased. SalUn [123] introduces the concept of “weight saliency” by focusing on specific model weights rather than the entire model. Dark Miner [124] follows a recurring three-stage process of mining, verifying, and circumventing. It greedily mines embeddings with the highest generation probabilities for unsafe concepts and effectively reduces the generation of unsafe content by targeting and mitigating these high-risk embeddings. EraseDiff [125] frames the unlearning task as a constrained optimization problem. It designs an inner optimization and an outer optimization. The former focuses on erasing undesirable influence by optimizing the models to fail to generate meaningful images corresponding to unsafe concepts. The latter preserves model performance by optimizing with the original diffusion loss function. Other works encourage [126, 127, 128] the forgetting of undesirable information in diffusion models by aligning the conditional scores of unsafe concepts with those of safe ones. SFD [126] integrates a score-based loss into the score distillation objective of a pre-trained diffusion model. SDD [127] applies self-distillation to the diffusion model, guiding the noise estimate conditioned on the target concept to align with the unconditional estimate. SafeGen [128] removes unsafe representations by projecting them into images with thick mosaics.
Unlike the above methods which design the noise prediction loss, some works [129, 130, 131] modify the attention weights $W$ to induce targeted changes in the keys and values of erased concepts $c_{i}$ and minimize changes of preserved concepts $c_{j}$. Specifically, these approaches aim to find weights such that the output for each input $c_{i}$ maps to safe values $v_{i}^{*}$ rather than the original $W^{\mathrm{old}}c_{i}$, while preserving the outputs for the inputs $c_{j}$ as $W^{\mathrm{old}}c_{j}$. A formal objective function can be constructed as follows:
$\min_{W}\sum_{c_{i}\in E}||Wc_{i}-v_{i}^{*}||_{2}^{2}+\sum_{c_{j}\in P}||Wc_{j}-W^{\mathrm{old}}c_{j}||_{2}^{2}.$ (12)
Notably, the objective function in Equation 12 has a closed-form solution for the updated weights:
$W=\text{\footnotesize$\left(\sum_{c_{i}\in E}v_{i}^{*}c_{i}^{T}+\sum_{c_{j}\in P}W^{\mathrm{old}}c_{j}c_{j}^{T}\right)\left(\sum_{c_{i}\in E}c_{i}c_{i}^{T}+\sum_{c_{j}\in P}c_{j}c_{j}^{T}\right)^{-1}$}.$ (13)
Specifically, UCE [129] uses the closed-form solution to optimize the model’s attention weights, selectively erasing the target concepts while minimizing alterations to the preserved concepts. MACE [130] expands the erasure capacity to handle up to 100 concepts by combining the closed-form cross-attention refinement with LoRA fine-tuning. RECE [131] employs closed-form parameter editing combined with adversarial learning schemes to achieve reliable and efficient concept erasing.
Reinforcement learning is also introduced to erase unsafe concepts [132, 133]. ShieldDiff [132] tackles unsafe content removal by fine-tuning a pre-trained diffusion model through reinforcement learning. It employs a customized reward function that combines the CLIP model with nudity-specific rewards to filter out the nudity content. DUO [133] removes unsafe content from T2I models through preference optimization with paired image data.
There are also novel approaches that do not fall into the above categories. For example, Forget-Me-Not [134] introduces attention Re-steering which fine-tunes the UNet component to minimize the attention maps corresponding to the target concepts. UC [135] proposes a domain correction framework using adversarial training to align sensitive and anchor concepts in diffusion models, along with gradient surgery to preserve model performance by mitigating conflicts between unlearning and relearning gradients.
@@
$\bullet$ Image + Text → Text (IT2T) Models. Some works [136, 137] apply safety steering vectors to calibrated activation at inference time, aiming to address harmful intents while preserving the model’s performance on benign inputs. Specifically, InferAligner [136] extracts steering vectors by calculating the activation difference between harmful and harmless prompts, and employs a guidance gate to selectively control activation shifts in specific transformer layers. Based on InferAligner, ASTRA [137] identifies visual tokens from adversarial images that are most strongly associated with jailbreaks and uses these tokens to construct steering vectors.
$\bullet$ Text → Image (T2I) Models. EIUP [138] mitigates unsafe content by re-weighting the attention map of the target token, guided by an introduced target unsafe concept. Specifically, it integrates image latent variables with the latent embedding of the erasure prompt to generate a corresponding target unsafe attention map. During the attention feature adjustment stage, the target unsafe attention map is merged with the attention map of the original prompt. Subsequently, the target unsafe attention map is reweighted to suppress unsafe features. SLD [139] mitigates inappropriate content generation by combining text conditioning with classifier-free guidance. It edits images during inference without fine-tuning, using unsafe prompts to guide generation in the opposite direction.
$\bullet$ Text → Video (T2V) Models. SAFREE [140] identifies unsafe tokens and projects them into a space orthogonal to the unsafe concept subspace, while retaining their representations within the original input space. To balance toxicity filtering and the preservation of safe concepts, it employs a self-validating filtering mechanism, dynamically adjusting denoising steps and using adaptive re-attention within the diffusion latent space.
Pruning-based Defense. Unsafe content often originates from concept neurons, which are essential for generating specific concepts. Pruning model parameters to deactivate unwanted concept neurons offers an effective strategy for eliminating harmful content.
$\bullet$ Text → Image (T2I) Models. Some works [141, 142] selectively remove critical parameters linked to the undesired concepts. Specifically, P-ESD [141] enhances concept-erasing techniques by identifying concept-correlated neurons that are sensitive to adversarial prompts. ConceptPrune [142] identifies skilled neurons in feed-forward layers responsible for undesirable concepts in diffusion models.